Jill is 8 kilometers away from Joe. Both began to walk toward each other at the same time, Jill walks at 1. kilometers and hour they meet in two hours, how fast is Joe walking
Answer by ptaylor(2048)

Distance(d) = rate(r) times time(t) or d=rt and r=d/t
Let r= Joe's rate (how fast he's walking)
In 2 hours Jill walks 2 kilometers (d=1*2) that leaves 6 kilometers (8-2) for Joe to walk in 2 hours
r=6/2=3 kilometers/hour-----------------------------------Joe's speed
CK
1 kilometer/hour times 2 hours plus 3 kilometers/hour times 2 hours =8 kilometers
8 kilometers=8 kilometers
Both together walked 8 kilometers
Hope this helps-----ptaylor